OpenStudy (anonymous):

Using the given zero, find all other zeros of f(x). -2i is a zero of f(x) = x^4 - 5x^2 - 36 These are the answer choices: (a)2i, 6i, -6i (b)2i, 6, -6 (c)2i, 3, -3 (d)2i, 3i, -3i

OpenStudy (triciaal):

A zero of a function is the value of x when the function =0. When the function = 0. It means when divided by the factors the remainder =.0 Given -2i is a zero then factor (x + 2i)=o Complex numbers come in pairs so (x.- 2i) is also a factor

OpenStudy (anonymous):

Then what's the answer?

OpenStudy (triciaal):

Divide f(x) by the factors

OpenStudy (anonymous):

\[\frac{x^4-5x^2-36}{(x-2i)(x+2i)}\]

OpenStudy (triciaal):

Example if 2 is a factor of 24. Then to find the other factor divide 24 by 2 = 12. So 2 * 12 = 24 To get all the factors need to break down 12 as well For example 24 = 2*3*4

OpenStudy (anonymous):

what does the i stand for?

OpenStudy (triciaal):

The numerator can be factorized Quadratic. Then cancel the common term

OpenStudy (anonymous):

i is the square root of negative one

OpenStudy (triciaal):

I = square root of -1

OpenStudy (triciaal):

I squared = -1

OpenStudy (anonymous):

\[\frac{x^4-5x^2-36}{(x-2i)(x+2i)}=\frac{x^4-5x^2-36}{(x^2+4)}=\frac{x^4+(4x^2-9x^2)-36}{(x^2+4)}\] \[\frac{(x^4+4x^2)-(9x^2+36)}{(x^2+4)}=\frac{(x^4+4x^2)}{x^2+4}-\frac{(9x^2+36)}{(x^2+4)}\] \[x^2-9\] \[(x+3)(x-3)\]

OpenStudy (triciaal):

A negative number can be -1 times the number To get the square root it would be rt of -1 times the rt of the number Rt of -1 is denoted I.

OpenStudy (anonymous):

I can show you how to multiply (x-2i)(x+2i)

OpenStudy (anonymous):

wait, is it C?

OpenStudy (anonymous):

Maybe

OpenStudy (anonymous):

i think I got it

OpenStudy (anonymous):

how would i be able to find all the answers then?

OpenStudy (anonymous):

So a 4th order polynomial (x is raised to the 4th power) has 4 roots.

OpenStudy (anonymous):

They give you one root: (x+2i)

OpenStudy (anonymous):

An important piece of information is that if all of the coefficients are real, then the complex root (x-2i) MUST be accompanied by its complex conjugate, which is just (x-2i).

OpenStudy (anonymous):

Now you know 2 roots. You can combine the roots together to make something that doesn't have i in it.

OpenStudy (anonymous):

Wait I get it!

OpenStudy (anonymous):

It's C!

OpenStudy (anonymous):

(x-2i)(x+2i) =x(x+2i)-2i(x+2i) =x^2+(2*x*i)-(2*i*x)+2*2*i^2 =x^2+4i^2 =x^2-4 Remember i is the sqrt of -1